Present: Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying TangAnnouncements
Ongoing Systems Updates
- The State Auditor will be running a security scan on our network on Thursday, February 18 - between 11:30 - 1:30 PM. This may cause network slowdowns or crashes. They will be testing Windows '95 & '98 work stations. Don't plan on doing much during that time period. Scan results should be immediate.
- Caroline - Unix box - AIX 3.2 is not Y2K compliant. Plans were to migrate off of it by the end of the year - will start doing so now. The security hole with it is SEND mail. Most of our use with it is web-based. Suzy hopes to have it migrated by Thursday morning, if not the SEND mail will be shut off. Suzy has been trying to find a last-minute upgrade with no luck. IBM has announced that it will support RS6000 Linux - we may switch Caroline over to it.
- NT Spectre software has been received. It performs internal checks on passwords. Upon its first run, it found only half a dozen inappropriate passwords. Those have been set to expire. We will run the software every couple of weeks.
- Barcode scanners were received - distribution list has been issued and 12 more scanners are on order.
- Bill will alert the Liaisons to the Security Scan.
Software Projects Status
- We're currently waiting on cost quotes from Global & Digital. There was a problem with Nordex - they use 110 tools. Ashley requested a tool, they'll check on it. Wiring should come with a lifetime warranty.
- Upgrades planned for the MSL closet, Music AFA, Library West & Resource Services. May want to tackle these during Spring Break as it may cause network slowness, possible interruption.
- Ann informed Ashley that Ed Teague would want to review the AFA plan first.
- MSL will use 3COM equipment to add the fiber switch.
- When these projects are complete, all of the closets and users should be at CAT5 or above to handle 100M.
- Campus may pay the cost of the switch at the POP.
- Documentation for the wiring projects is in process, Ashley is trying to get it online and update it to include before & after.
- Winston is finishing up on Mellon changes.
- Ying is working on the IR database. She'll be meeting with Bobbie this week to discuss what needs to be kept or retracted. Bill would like to see any proposed changes.
- Brian has been working on LASSIE - using InterDev - can't seem to make the whole thing happen. He can get page display, but it's not pulling data. There are other alternatives - CGI interface, for one. InterDev is supposed to be an easier way.
- Compustat Yearly tapes - Sharon Wheeler at NERDC was unable to read the 3270 cartridges. Bill will get with Peter McKay to get in touch with the marketing representative. CD format would be easier to deal with.
- Ashley out Friday, February 19.
Last updated February 18, 2000
by Debra Harris